OpenStudy (anonymous):

Serena saved eighty percent of the money she made babysitting last week. If she made sixty-five dollars last week, how much did she save? Show your work.

jimthompson5910 (jim_thompson5910):

80% of 65 = (80/100)*65 = 0.8*65 = ??

OpenStudy (anonymous):

52?

jimthompson5910 (jim_thompson5910):

so that's how much she saved

OpenStudy (anonymous):

Thankss!

jimthompson5910 (jim_thompson5910):

yw